# Temporal Sepsis Modeling: a Fully Interpretable Relational Way

### v1

@@ Abstract
Sepsis remains one of the most complex and heterogeneous syndromes in intensive care, characterized by diverse physiological trajectories and variable responses to treatment. While deep learning models perform well in the early prediction of sepsis, they often lack interpretability and ignore latent patient sub-phenotypes. In this work, we propose a machine learning framework by opening up a new avenue for addressing this issue: a relational approach. Temporal data from electronic medical records (EMRs) are viewed as multivariate patient logs and represented in a relational data schema. Then, a propositionalisation technique (based on classic aggregation/selection functions from the field of relational data) is applied to construct interpretable features to “flatten” the data. Finally, the flattened data is classified using a selective naive Bayesian classifier. Experimental validation demonstrates the relevance of the suggested approach as well as its extreme interpretability. The interpretation is fourfold: univariate, global, local, and counterfactual.
I Introduction
@@
Sepsis [28] is defined as a life-threatening organ dysfunction caused by an aberrant dysregulated host response to infection, and progression to septic shock markedly increases the risk of multi-organ failure and death. Early identification and risk stratification using validated clinical criteria, such as increases in SOFA score, enable rapid implementation of evidence-based interventions—including prompt broad-spectrum antimicrobial therapy and targeted resuscitation—which are associated with improved outcomes. Despite these advances, sepsis remains a leading cause of morbidity and mortality in intensive care worldwide.
@@
Despite decades of clinical and computational research, early detection remains a major challenge due to the disease’s heterogeneity and temporal complexity. Traditional machine learning models treat sepsis as a binary event, neglecting the underlying patient-specific physiological patterns that drive its onset and progression. However, early detection of sepsis using machine learning algorithms remains an area of research that is still in an ongoing stage of development. Despite recent advances, existing models exhibit limited performance, with some failing to identify up to 67% of sepsis cases [38], which severely restricts their deployment in real-world clinical settings. To ensure that these tools can be used reliably across diverse patient populations, further studies focusing on their clinical implementation and evaluation under real-world conditions are required [10]. Moreover, validating the reproducibility and generalization capabilities of these models remains a major challenge [30]. This difficulty is attributable to the fact that most existing studies have mainly focused on populations from intensive care units (ICUs) [32], thereby limiting the applicability of these models to other hospital departments. In addition, several retrospective studies have demonstrated that the success of machine learning models for sepsis detection depends heavily on the integration of clinically relevant and informative variables [10]. Nevertheless, the development and evaluation of predictive models leveraging data from hospitalized patients across all clinical services remain limited. In this context, the MIMIC database provides a broad panel of clinical, biological, and physiological parameters, making it particularly well suited for studying and comparing early sepsis prediction approaches. Accordingly, MIMIC-III [13] has become a widely used benchmark dataset in the literature for the comparative evaluation of sepsis prediction models and constitutes the experimental framework adopted in this study.
@@
This paper is organized as follows. Section II reviews the main families of sepsis modeling. Section III details the relational representation of patient trajectories, the technique of propositionalisation by flattening, and variable selection, emphasizing the ability of this method to produce understandable characteristics. Section IV then presents the performance of the models, the quality of the aggregated variables, and the analysis of their univariate and local importance, while introducing the notion of counterfactual reasoning to explain and influence the model’s decisions. Finally, the last section summarizes the contributions of the proposed approach, highlighting its advantages in terms of explainability and flexibility.
II Context and Related Works
@@
II-A Time-series and deep learning approaches
@@
II-B Feature-based and window-aggregation approaches
@@
II-C MDL-based and Bayesian explainable models
@@
While relational feature extraction provides a systematic way to transform patient trajectories into a flat attribute-value table suitable for classical machine learning, the large number of potential aggregate features can easily lead to overfitting and redundant representations. To address these limitations, Minimum Description Length (MDL)–based methods offer a principled framework for both constructing and selecting informative features. By explicitly balancing model fit and complexity, MDL-based approaches generate compact, high-quality aggregates that capture temporal patterns while remaining suitable for interpretable probabilistic classifiers [5].
Explainability is further enhanced by selective Bayesian classifiers such as Fractional Naive Bayes [12], which aim to retain only informative and weakly dependent variables. In contrast to post-hoc explainability techniques applied to black-box models (e.g., SHAP [23]), these approaches provide intrinsic interpretability through probabilistic reasoning and explicit variable selection.
II-D Prediction horizons in early sepsis modeling
The prediction horizon, denoted $h$, defines the lead time before sepsis onset at which a model aims to anticipate the condition. In classical machine learning approaches based on aggregated features [6, 35], $h$ is typically fixed by the aggregation window: only measurements up to $h$ hours before onset are included in feature computation. This constrains the models to rely on recent observations and may limit long-range predictive power.
@@ II-D Prediction horizons in early sepsis modeling
Deep learning models such as LSTM, TCN, and MGP-RNN [11, 26] inherently handle sequences of arbitrary length and can be trained to predict multiple horizons simultaneously. For instance, models can be trained to output risk scores at $h=1,3,6$ hours before sepsis onset, allowing a flexible trade-off between early detection and predictive accuracy. Multi-horizon prediction is particularly advantageous in ICU settings, where early warnings must be balanced against false alarms.
@@
For relational and MDL-based methods [5, 12], the prediction horizon is explicitly controlled through the number of temporal observations $p$ included in the secondary table. Each patient trajectory is flattened over $p$ time steps corresponding to the desired horizon $h$, enabling consistent learning of aggregate variables that encode temporal dynamics relevant for early prediction. This approach allows the use of interpretable probabilistic classifiers while maintaining horizon-specific information. Overall, while deep learning methods naturally support multi-horizon prediction, feature-based and relational methods achieve horizon flexibility by careful window selection and aggregation, highlighting a key design choice in early sepsis modeling.
@@
To address the interpretability limitations of existing methods and lack of principled complexity control we represent patient trajectories as a root table linked to a secondary table of temporal observations. MDL-based aggregation, combined with supervised discretization and selective Bayesian classification, yields compact, interpretable features suitable for varying prediction horizons.
Our methodology details (i) relational data representation, (ii) feature extraction and flattening, and (iii) MDL-driven variable selection and Bayesian classification.
III Modeling: From temporal Data to Flat Data
III-A Dataset used
@@
III-B Viewing the data as relational data
The dataset described in the previous section may be viewed as a relational dataset (see Figure 1) with a star schema. In this paradigm, the root table contains two columns: the first one is the ‘Id’ of the patient (a reference to the patient), and the second one contains the class to predict (healthy or sick). In the secondary table, the number of columns corresponds to the $d$ ($d=34$) explanatory variables described in the previous section plus the column ‘Id’ to enable table joins. The number of lines of the secondary table is $pN$, where $p$ is the number of (logs or time stamps) of each patient used to do the prediction, which depends on the horizon for which one would like the prediction, as described just below.
Figure 1: Time series relational data flattening
This is represented in the Figure 2. In the initial dataset, each patient is represented by a measure (a vector) of 34 variables, each one measured 12 times ($t\in\{t_{-6},...,t_{+5}\}$) every hour, so at different time stamps. In the dataset used all the patients become “sick” a $t=t_{0}$. Therefore to predict the nature of the patient one hour before $t=t_{0}$ one may used a maximum of 6 measures ($h=1$ and $p=6$) and for a prediction 3 hours before $t=t_{0}$ one may used a maximum of 4 measures ($h=3$ and $p=4$) etc. Indeed, the number of lines, $p$, to keep in the secondary table depends on the value of $h$.
@@ III-B Viewing the data as relational data
Note 1: using the same value for $p$ and $h$ (see Figure 1) for all patients does not in any way reduce the scope of the analysis. The proposed approach works in exactly the same way if $p$ (with $h$ set to the same value for fairness to patients) were to vary from patient to patient (if the latter had more than 4 measurements in the past before $t_{-3}$).
Note 2: The proposed approach could also be multi-modal. Indeed, it would suffice to have one (or more) other table(s) containing alternative representations of patients and to apply the same process.
III-C Flatening the data
Machine learning models typically require input data in a tabular, flat format where each row represents an instance and each column represents a feature. Relational data stored across multiple tables needs to be combined into a single table to be used effectively for training classifiers. This action is also named propositionalization.
The propositionalization approaches consist of learning a model from relational data by flattening the original relational data, which are stored in several linked tables (similarly to databases) [16]. These approaches come from the field of Relational Data Mining [9] and are not usually applied to time series. More precisely, relational data contains at least one root table where each row represents a statistical individual (e.g. ‘id‘ of the patients) and another secondary table containing detailed records (eg. the measurements realized of patients which are represented by each row of the secondary table). The specificity of relational data is to involve one-to-many relationships between the tables (eg. a patient underwent several identical medical examinations over time). The propositionalization problem consists of transforming the relational data into a single attribute-value dataset in order to use regular machine learning methods.
@@ III-C Flatening the data
Our relational data (see Fig. 1) consists of the root table, which contains $N$ instances, characterized by two variables: the patient identifiers and the target variable (i.e. class values). The secondary table contains $pN$ detailed medical examination, described by 36 variables already described in Section III-A. The number of aggregate variables to be constructed ($Q$) is the only user parameter. Our relational data is transformed into a regular attribute-value dataset by applying the MDL based propositionalization approach.
III-D Variable selection and learned classifier
@@
III-D1 Filtering uninformative variables
The filtering of the aggregate variables is based on previously developed supervised discretization ([4]) and grouping ([3]) methods. This Bayesian approach turns the learning task into a model selection problem. A prior distribution is defined on the model space that exploits the hierarchy of their parameters. In practice, this approach reaches a good trade-off between robust and accurate models. The prior favors simple models with few intervals, and the likelihood favors models that fit the data regardless of their complexity. The aggregate variables are evaluated, one by one, using a specifically designed MDL optimization criterion ([5]). The complexity of the aggregate variables is taken into account by adding a construction cost in the prior. This criterion can be interpreted as a coding length according to information theory. Compression Gain (CG) compares the coding length of the learned model with the empty model that includes a single interval. CG measures the ability of the learned models to compress the training data, despite the additional construction cost. Only the $R$ variables with a positive CG are retained ($R\leq Q$)
III-D2 Finding the most informative subset of variables:
@@ III-D3 Learning classifier
$C$ is the number of class values to be predicted (in this paper $C=2$), $V$ the input variables and $W$ the weights on the variable coming from the process described in [12]. This simple and efficient classifier makes the assumption that the distributions $P(V_{j}=x_{jk}|C_{z})$ are independent. In practice, these conditional distributions are estimated in a frequentist way, by using the previously learned univariate discretization ([4]) and grouping ([3]) models. The denominator of Eq. 1 normalizes the estimated probability by making a sum of the numerator term over all the class values. At the end, the predicted class value given a particular $x_{k}$ is the one that maximizes the conditional probabilities $P(C_{z}|x_{k})$.
Note: Other classifiers than the SNB here used could be employed with the S, R or Q variables found with the process described above. In this paper for the sake of the complete line with a MDL principle (and also since the results are quite good) we did not test this possibility but interested readers may try. All the results are fully reproducible with the Github where the dataset and the used code are available.
IV Results
IV-A Classifier performances
The average results of the classifier on a 10-fold cross-validation are presented in Table I in terms of both accuracy and AUC, compared to the value of the number of $Q$ variables to be constructed (as described in the previous section).
@@ IV-A Classifier performances
Table I: Caption
Obviously the performances are very good. If a tradeoff between the number of variables incorporated in the model ($R$) and the AUC in test is required, then the value QQ=1000, RR=872 and RR=98 could be considered. The 98 aggregate variables to be constructed are sufficient to obtain good performance and a frugal model. But it is up to the final user since the largest model has only 272 input variables. The next section will present some of these aggregate variables.
IV-B Example of aggregates elaborated
The Table II, column 1, gives the $S$ 20 most important aggregated variables found by the approach used [5] here. A nice property is their straightforward meaning. On this dataset they have only two operands, even if the method in [5] may produce more complex “rules”. Each of them is easy to read and to understand even for a non-data scientist (or research scientist) and moreover in terms readable by medical practitioners. We also give for the first one a color code where the blue parts are the rules and the orange parts the variables present in the secondary table (to ease the reading). It is worth noting that the clinical interpretation of some of these rules still needs to be validated using longitudinal data collected between 2022 and 2025 at XXXX [^1] The name of the hospital is not here given for anonymity reason during submission of this paper. Hospital.
Name Level Importance
@@ IV-B Example of aggregates elaborated
Table II: The $S$ 20 most important aggregated variables
IV-C Univariate variable importance
As described in Section III-D the Khiops classifier is trained in two steps. The first one is to evaluate the univariate predictive information contained in each aggregated variable. For that, a supervised discretization [4] is realized for numerical one and a supervised grouping method [3] is performed on categorical variables. In the dataset used here all the variables are numerical (except the ’Gender’ and ‘Hour’ variables redrawn at the beginning of the study). One output of this preprocessing is an information of the univariate predictive information, named “level”, of each variable which is in the range $[0:1]$ (0 variable no informative, 1 variable 100 % informative). These values are given in the second column of the Table I.
Since each aggregated variable is discretized, it is also possible to have a look at the information carried by the intervals found. For that the Khiops library provides an interactive results visualization tool, called Khiops Visualization. We present in Figure 3 a copy of a screenshot of this tool for the most important aggregate: $V_{1}$=“Min(EtCO2) where Calcium >> 8.075”.

### v2

@@ Abstract
Sepsis remains one of the most complex and heterogeneous syndromes in intensive care, characterized by diverse physiological trajectories and variable responses to treatment. While deep learning models perform well in the early prediction of sepsis, they often lack interpretability and ignore latent patient sub-phenotypes. In this work, we propose a machine learning framework by opening up a new avenue for addressing this issue: a relational approach. Temporal data from electronic medical records are viewed as multivariate patient logs and represented in a relational data schema. Then, a propositionalisation technique (based on classic aggregation/selection functions from the field of relational data) is applied to construct interpretable features to “flatten” the data. Finally, the flattened data is classified using a selective naive Bayesian classifier. Experimental validation demonstrates the relevance of the suggested approach as well as its extreme interpretability. The interpretation is fourfold: univariate, global, local, and counterfactual.
1 Introduction
@@
Sepsis [26] is defined as a life-threatening organ dysfunction caused by an aberrant dysregulated host response to infection, and progression to septic shock markedly increases the risk of multi-organ failure and death. Early identification and risk stratification using validated clinical criteria, such as increases in SOFA score, enable rapid implementation of evidence-based interventions, including prompt broad-spectrum antimicrobial therapy and targeted resuscitation, which are associated with improved outcomes. Despite these advances, sepsis remains a leading cause of morbidity and mortality in intensive care worldwide.
@@
Despite decades of clinical and computational research, early detection remains a major challenge due to the disease’s heterogeneity and temporal complexity. Traditional machine learning models treat sepsis as a binary event, neglecting the underlying patient-specific physiological patterns that drive its onset and progression. However, early detection of sepsis using machine learning algorithms remains an area of research that is still in an ongoing stage of development. Despite recent advances, existing models exhibit limited performance, with some failing to identify up to 67% of sepsis cases [35], which severely restricts their deployment in real-world clinical settings. To ensure that these tools can be used reliably across diverse patient populations, further studies focusing on their clinical implementation and evaluation under real-world conditions are required [9]. Moreover, validating the reproducibility and generalization capabilities of these models remains a major challenge [28]. This difficulty is attributable to the fact that most existing studies have mainly focused on populations from intensive care units (ICUs) [30], thereby limiting the applicability of these models to other hospital departments. In addition, several retrospective studies have demonstrated that the success of machine learning models for sepsis detection depends heavily on the integration of clinically relevant and informative variables [9]. Nevertheless, the development and evaluation of predictive models leveraging data from hospitalized patients across all clinical services remain limited. In this context, the MIMIC database provides a broad panel of clinical, biological, and physiological parameters, making it particularly well-suited for studying and comparing early sepsis prediction approaches. Accordingly, MIMIC-III [12] has become a widely used benchmark dataset in the literature for the comparative evaluation of sepsis prediction models and constitutes the experimental framework adopted in this study.
@@
This paper is organized as follows. Section 2 reviews the main families of sepsis modeling. Section 3 details the relational representation of patient trajectories, the technique of propositionalisation by flattening, and variable selection, emphasizing the ability of this method to produce understandable characteristics. Section 4 then presents the models’ performances, the quality of the aggregated variables, and the analysis of their univariate and local importance, while introducing counterfactual reasoning to explain and influence the model’s decisions. Finally, the last section summarizes the contributions of the proposed approach, highlighting its advantages in terms of explainability and flexibility.
2 Context and Related Works
@@
2.1 Time-series and deep learning approaches
@@
2.2 Feature-based and window-aggregation approaches
@@
2.3 MDL-based and Bayesian explainable models
@@
While relational feature extraction provides a systematic way to transform patient trajectories into a flat attribute-value table suitable for classical machine learning, the large number of potential aggregate features can easily lead to overfitting and redundant representations. To address these limitations, Minimum Description Length (MDL)–based methods offer a principled framework for both constructing and selecting informative features. By explicitly balancing model fit and complexity, MDL-based approaches generate compact, high-quality aggregates that capture temporal patterns while remaining suitable for interpretable probabilistic classifiers [5]. Explainability is further enhanced by selective Bayesian classifiers such as Fractional Naive Bayes [11], which aim to retain only informative and weakly dependent variables. In contrast to post-hoc explainability techniques applied to black-box models (e.g., SHAP [21]), these approaches provide intrinsic interpretability through probabilistic reasoning and explicit variable selection.
2.4 Prediction horizons in early sepsis modeling
@@
The prediction horizon, denoted $h$, defines the lead time before sepsis onset at which a model aims to anticipate the condition. In classical machine learning approaches based on aggregated features [6, 33], $h$ is typically fixed by the aggregation window: only measurements up to $h$ hours before onset are included in feature computation. This constrains the models to rely on recent observations and may limit long-range predictive power. Deep learning models such as LSTM, TCN, and MGP-RNN [24] inherently handle sequences of arbitrary length and can be trained to predict multiple horizons simultaneously. For instance, models can be trained to output risk scores at $h=1,3,6$ hours before sepsis onset, allowing a flexible trade-off between early detection and predictive accuracy. Multi-horizon prediction is particularly advantageous in ICU settings, where early warnings must be balanced against false alarms.
@@
For relational and MDL-based methods [11], the prediction horizon is explicitly controlled through the number of temporal observations $p$ included in the secondary table. Each patient trajectory is flattened over $p$ time steps corresponding to the desired horizon $h$, enabling consistent learning of aggregate variables that encode temporal dynamics relevant for early prediction. This approach allows the use of interpretable probabilistic classifiers while maintaining horizon-specific information. Overall, while deep learning methods naturally support multi-horizon prediction, feature-based and relational methods achieve horizon flexibility by careful window selection and aggregation, highlighting a key design choice in early sepsis modeling.
@@
To address the interpretability limitations of existing methods and the lack of principled complexity control, we represent patient trajectories as a root table linked to a secondary table of temporal observations. MDL-based aggregation, combined with supervised discretization and selective Bayesian classification, yields compact, interpretable features suitable for varying prediction horizons. Our methodology details (i) relational data representation, (ii) feature extraction and flattening, and (iii) MDL-driven variable selection and Bayesian classification.
3 Modeling: From temporal Data to Flat Data
3.1 Dataset used
@@
3.2 Viewing the data as relational data
The dataset described in the previous section may be viewed as a relational dataset (see Figure 1) with a star schema. In this paradigm, the root table contains two columns: the first one is the ‘Id’ of the patient (a reference to the patient), and the second one contains the class to predict (healthy or sick). In the secondary table, the number of columns corresponds to the $d$ ($d=34$) explanatory variables described in the previous section, plus the column ‘Id’ to enable table joins. The number of lines of the secondary table is $pN$, where $p$ is the number of (logs or time stamps) of each patient used to do the prediction, which depends on the horizon for which one would like the prediction, as described just below.
Figure 1: A Time series relational data flattening
@@ 3.2 Viewing the data as relational data
This is represented in the Figure 2. In the initial dataset, each patient is represented by a measure (a vector) of 34 variables, each one measured 12 times ($t\in\{t_{-6},...,t_{+5}\}$) every hour, so at different time stamps. In the dataset used all the patients become “sick” a $t=t_{0}$. Therefore, to predict the nature of the patient one hour before $t=t_{0}$ one may use a maximum of 6 measures ($h=1$ and $p=6$) and for a prediction 3 hours before $t=t_{0}$ one may used a maximum of 4 measures ($h=3$ and $p=4$) etc. Indeed, the number of lines, $p$, to keep in the secondary table depends on the value of $h$.
Notes: (i) using the same value for $p$ and $h$ (see Figure 1) for all patients does not in any way reduce the scope of the analysis. The proposed approach works in exactly the same way if $p$ (with $h$ set to the same value for fairness to patients) were to vary from patient to patient (if the latter had more than 4 measurements in the past before $t_{-3}$). (ii) The proposed approach could also be multi-modal. Indeed, it would suffice to have one (or more) other table(s) containing alternative representations of patients and to apply the same process.
3.3 Flatening the data
Machine learning models typically require input data in a tabular, flat format where each row represents an instance and each column represents a feature. Relational data stored across multiple tables needs to be combined into a single table to be used effectively for training classifiers. This action is also named propositionalization. The propositionalization approaches consist of learning a model from relational data by flattening the original relational data, which are stored in several linked tables (similarly to databases) [15]. These approaches come from the field of Relational Data Mining [8] and are not usually applied to time series. More precisely, relational data contains at least one root table where each row represents a statistical individual (e.g. ‘id‘ of the patients) and another secondary table containing detailed records (eg. the measurements realized of patients which are represented by each row of the secondary table). The specificity of relational data involves one-to-many relationships between the tables (eg. a patient underwent several identical medical examinations over time). The propositionalization problem consists of transforming the relational data into a single attribute-value dataset in order to use regular machine learning methods.
@@ 3.3 Flatening the data
Our relational data (see Fig. 1) consists of the root table, which contains $N$ instances, characterized by two variables: the patient identifiers and the target variable (i.e., class values). The secondary table contains $pN$ detailed medical examination, described by 36 variables already described in Section 3.1. The number of aggregate variables to be constructed ($Q$) is the only user parameter. Our relational data is transformed into a regular attribute-value dataset by applying the MDL-based propositionalization approach.
3.4 Variable selection and learned classifier
@@
3.4.1 Filtering uninformative variables:
The filtering of the aggregate variables is based on previously developed supervised discretization ([4]) and grouping ([3]) methods. This Bayesian approach turns the learning task into a model selection problem. A prior distribution is defined on the model space that exploits the hierarchy of its parameters. In practice, this approach reaches a good trade-off between robust and accurate models. The prior favors simple models with few intervals, and the likelihood favors models that fit the data regardless of their complexity. The aggregate variables are evaluated, one by one, using a specifically designed MDL optimization criterion ([5]). The complexity of the aggregate variables is taken into account by adding a construction cost in the prior. This criterion can be interpreted as a coding length according to information theory. Compression Gain (CG) compares the coding length of the learned model with the empty model that includes a single interval. CG measures the ability of the learned models to compress the training data, despite the additional construction cost. Only the $R$ variables with a positive CG are retained ($R\leq Q$)
3.4.2 Finding the most informative subset of variables:
@@ 3.4.3 Learning classifier:
$C$ is the number of class values to be predicted (in this paper $C=2$), $V$ the input variables, and $W$ the weights on the variable coming from the process described in [11]. This simple and efficient classifier makes the assumption that the distributions $P(V_{j}=x_{jk}|C_{z})$ are independent. In practice, these conditional distributions are estimated in a frequentist way, by using the previously learned univariate discretization ([4]) and grouping ([3]) models. The denominator of Eq. 1 normalizes the estimated probability by making a sum of the numerator term over all the class values. At the end, the predicted class value given a particular $x_{k}$ is the one that maximizes the conditional probabilities $P(C_{z}|x_{k})$.
Note: Other classifiers than the SNB here used could be employed with the S, R or Q variables found with the process described above. See the appendix in the GitHub.
4 Results
4.1 Classifier performances
The average results of the classifier on a 10-fold cross-validation are presented in Table 3 in terms of both accuracy and AUC, compared to the value of the number of $Q$ variables to be constructed (as described in the previous section).
Obviously, the performances are very good. If a tradeoff between the number of variables incorporated in the model ($R$) and the AUC in test is required, then the values QQ=1000, RR=872 and RR=98 could be considered.
@@ 4.1 Classifier performances
Table 1: AUC Results versus the number of Q variables to be constructed.
The 98 aggregate variables to be constructed are sufficient to obtain good performance and a frugal model. But it is up to the final user since the largest model has only 272 input variables. The next section will present some of these aggregate variables.
4.2 Example of aggregates elaborated
The Table 3, column 1, gives the $S$ 20 most important aggregated variables found by the approach used [5] here. A nice property is their straightforward meaning.
On this dataset they only have two operands, even if the method in [5] may produce more complex “rules”. Each of them is easy to read and to understand even for a non-data scientist (or research scientist) and moreover, in terms readable by medical practitioners. We also give for the first one a color code where the blue parts are the rules and the orange parts the variables present in the secondary table (to ease the reading). It is worth noting that the clinical interpretation of some of these rules still needs to be validated using longitudinal data collected between 2022 and 2025 at XXXX [^1] The name of the hospital is not given here for anonymity reasons during submission of this paper. Hospital.
Name Level Imp
@@ 4.2 Example of aggregates elaborated
Figure 3: The $S$ 20 most important aggregated variables
4.3 Univariate variable importance
As described in Section 3.4 the Khiops classifier is trained in two steps. The first one is to evaluate the univariate predictive information contained in each aggregated variable. For that, a supervised discretization [4] is realized for numerical one and a supervised grouping method [3] is performed on categorical variables. In the dataset used here all the variables are numerical (except the ’Gender’ and ‘Hour’ variables redrawn at the beginning of the study). One output of this preprocessing is an information of the univariate predictive information, named “level”, of each variable, which is in the range $[0:1]$ (0 variable no informative, 1 variable 100 % informative). These values are given in the second column of the Table 3. Since each aggregated variable is discretized, it is also possible to have a look at the information carried by the intervals found. For that the Khiops library provides an interactive results visualization tool, called Khiops Visualization. We present in Figure 4 a copy of a screenshot of this tool for the most important aggregate: $V_{1}$=“Min(EtCO2) where Calcium >> 8.075”. The analysis of the discretization allows us to understand why this aggregate is informative. Small values of $V_{1}$ produce intervals relatively pure of Sepsis=1 while the intervals between 26 and 36 are relatively pure of Sepsis=0. The reading of $P(X|C)$ per interval, and for each variable, is informative to understand the result of the univariate classifications.